Write the equation of a line that passes through the points (-3,7) and (3,1).
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Write the equation of a line that passes through the points (-3,7) and (3,1).
-----
slope = m = (1-7)/(3--3) = -6/6 = -1
------------------
Form:: y = mx + b
Solve for "b"::
1 = -1*3 + b
b = 4
-----
Equation:
y = -x+4
-----------
Cheers,
Stan H.
---------------

Answer by solver91311(24713)

..

Use the Two-Point Form:



where and are the coordinates of the given points.

Plug in your numbers and then simplify.

John

My calculator said it, I believe it, that settles it